# On the applicability of Kolmogorov’s theory of probability

### v2

@@ Abstract
It is a common view that von Neumann laid the foundations of a “quantum probability theory” with his axiomatization of quantum mechanics (QM). As such, it is regarded a generalization of the “classical probability theory” due to Kolmogorov. Outside of quantum physics, however, Kolmogorov’s axioms enjoy universal applicability. This raises the question of whether quantum physics indeed requires such a generalization of our conception of probability or if von Neumann’s axiomatization of QM was contingent on the absence of a general theory of probability in the 1920s.
In this work I argue in favor of the latter position. In particular, I show that for non-relativistic $N$-body quantum systems subject to a time-independent scalar potential, it is possible to construct a mathematically rigorous theory based on Kolmogorov’s axioms and physically natural random variables, which reproduces central predictions of QM. The respective theories are distinct, so that an empirical comparison may be possible. Moreover, the approach can in principle be adapted to other classes of quantum-mechanical models.
Part II of this series will address the projection postulate and the question of measurement in this approach.
@@ 1 Introduction
For almost a century now, the theory of quantum mechanics (QM) has stood firm at the core of our modern understanding of the physical microcosm. Still, as evinced by the multitude of interpretations in existence today [freirejuniorOxfordHandbookHistory2022], the theory continues to mystify even experts in the subject. This state of affairs has given QM the idiosyncratic status of what one could call a well-established theory with contested foundations [freirejuniorQuantumDissidentsRebuilding2015, bacciagaluppiQuantumTheoryCrossroads2009].
@@
So far the following aspect of this problem has only received comparatively little attention [szaboCriticalReflectionsQuantum2001, accardiTopicsQuantumProbability1981, accardiNonKolmogorovProbability2009, fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952, nelsonDerivationSchrodingerEquation1966, redeiQuantumProbabilityTheory2007, streaterClassicalQuantumProbability2000, reddigerMadelungPictureFoundation2017, reddigerProbabilisticFoundationNonrelativistic2022, reddigerMathematicalTheoryMadelung2023]: across the empirical sciences, QM and its descendants are the only scientific theories that are not based on the mathematical theory of probability, as developed by Kolmogorov in 1933 [kolmogoroffGrundbegriffeWahrscheinlichkeitsrechnung1933]. Since von Neumann’s axiomatization of QM [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927, vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955] predates Kolmogorov’s work, it therefore stands to reason, whether quantum physics indeed requires a different notion of probability or if von Neumann’s axiomatization of QM was contingent on the absence of a general, mathematical theory of probability in the 1920s.
@@
In this work I argue in favor of the latter position. In particular, I show that for non-relativistic $N$-body quantum systems subject to a time-independent scalar potential, it is possible to construct a theory based on Kolmogorov’s axioms and physically natural random variables, which reproduces central predictions of QM (Lem. 1 and Thm. 1). The respective theories are distinct, so that their predictions can in principle be compared empirically.
@@
The main line of reasoning is as follows: First and with a historical focus, I elaborate on the general relationship of mathematical probability theory and QM (Sec. 2). In Sec. 3, a probabilistic view on the double slit experiment and quantum theory as a whole is provided, in line with the ensemble interpretation of QM [ballentineStatisticalInterpretationQuantum1970, pechenkinStatisticalEnsembleInterpretation2022]. This physically motivates the implementation of Kolmogorov’s axioms in quantum theory via the Born rule for position (Sec. 4, in particular Lem. 1). Corresponding observables can be obtained from Madelung’s reformulation of the Schrödinger equation (Sec. 5), in line with [reddigerMadelungPictureFoundation2017]. While this suggests that the theory can be erected upon the Madelung equations in conjunction with the Born rule, a recent review [reddigerMathematicalTheoryMadelung2023] showed that the mathematical theory of the Madelung equations is still underdeveloped. Hence and at least for the time being, another approach is needed. Instead, one may develop a “hybrid theory”, in which the dynamics is borrowed from QM and the basic quantities are rigorously defined in the context of said dynamics. For the aforementioned class of quantum systems this is done in Sec. 6. In result, an alternative theory based on Kolmogorov’s axioms is indeed obtained (Thm. 1).
@@ 2 Probability and quantum mechanics
The two modern mathematical theories of probability were developed by the Hungarian mathematician John von Neumann (1903–1957) [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927, vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955] and the Russian mathematician Andrej N. Kolmogorov (1903–1987) [kolmogoroffGrundbegriffeWahrscheinlichkeitsrechnung1933]. That today there are two such theories, points at a potential discrepancy in the modern foundations of probability theory.
@@
Historically, it is known that both researchers knew each other and followed each other’s work [monastyrskiiJohnNeumann2004]. Still, there does not seem to be any documented evidence of a discussion among the two about this particular topic. [^1] Neither an explicit search nor a consultation of [vonplatoCreatingModernProbability1994] yielded any result. Parthasarathy [parthasarathyObituaryAndreiNikolaevich1988] writes that Kolmogorov showed little interest in QM, adding that “it is a blessing to have enough ground left for future mathematicians to till.”
@@
It would go beyond the scope of this article to provide a direct comparison of the respective theories. Here I only note that in the contemporary literature this conflict is resolved by showing that von Neumann’s “quantum probability theory” may be viewed as a generalization of “classical probability theory,” as discovered by Kolmogorov [accardiTopicsQuantumProbability1981, streaterClassicalQuantumProbability2000, redeiQuantumProbabilityTheory2007, szaboCriticalReflectionsQuantum2001].
@@
While this accords with the general philosophy that QM is to be viewed as a generalization of “classical mechanics”, the particular historical context of the development of the two theories may be put forward as an argument against this conventional point of view.
@@
Mostly guided by the ideas of Bohr and Heisenberg, von Neumann developed his theory of probability with the explicit goal of axiomatizing QM. That – and not the development of a probability theory – was his aim. Indeed, in his most important article on the topic [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927] von Neumann noted that he was merely formalizing “the common calculus of probability.” Clearly, he was not aware that one day his notion of probability would be in conflict with what is generally associated with that term.
@@
Kolmogorov, on the other hand, took a direct approach to the question via measure theory. His fundamental work [kolmogoroffGrundbegriffeWahrscheinlichkeitsrechnung1933] was published in 1933. This is one year after von Neumann’s famous book on the mathematical foundations of QM [vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955] was published and a total of six years after von Neumann’s article on the topic [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927] appeared.
@@
One should add to this, that today Kolmogorov’s theory of probability is successfully employed in virtually all of the sciences—except quantum theory. Whereas von Neumann’s axiomatization is employed in quantum theory only. [^2] In [accardiNonKolmogorovProbability2009] supposed counterexamples are provided. Though it may be appropriate to discuss them point-by-point, I shall leave the reader with the remark that I find none of the given examples convincing.
Therefore, the hypotheses I put forward here are that, one, von Neumann’s axiomatization was contingent on the absence of a general, mathematical theory of probability at the time and, two, that the correct theory of probability is indeed the one discovered by Kolmogorov.
If this is the case, then contemporary quantum theory, both non-relativistic and relativistic, requires a revision with regards to the underlying probability theory.
In order to put these hypotheses to scientific scrutiny, the question of whether an empirically successful quantum theory can be constructed on the basis of Kolmogorov’s axioms is to be investigated. In what follows I shall show, that this may indeed be possible.
@@ 3 Physical considerations
Before addressing any mathematical details, I shall provide a brief introduction to the central physical ideas.
For this purpose, I begin with a discussion of the double slit experiment with material particles. [^3] I do not claim, that the discussion here applies to electromagnetic radiation as well. See Planck’s remark in [einsteinUeberEntwicklungUnserer1969]. Its character as an “examplar” in the sense of Kuhn [kuhnStructureScientificRevolutions2012] was also recognized by Feynman, who wrote that “it has in it the heart of quantum mechanics” and that “[i]n reality, it contains the only mystery” [feynmanFeynmanLecturesPhysics1963].
For the sake of argument, we shall assume that in the experiment only a single electron is released at a time, so that one can view the runs as mutually independent. Then, if an electron impinges on the detector screen, it creates a single, pointlike impact. Yet, as one repeats the experiment a large number of times, an interference pattern begins to emerge.
In the literature one commonly encounters the claim, that this pattern cannot result from particles in the usual sense of the word [feynmanFeynmanLecturesPhysics1963]. The argument is that the pattern is not the same as the conjunction of the two patterns one obtains with one of the slits closed. That is, if $\sigma_{\text{double}}$ is the ($2$-dimensional, time-independent) probability density of the impact locations on the detector screen for the double slit experiment and $\sigma_{i}$ with $i\in\{1,2\}$ are the probability densities for the experiment with one of the slits closed, then we have
@@
Indeed, if the presence of the other slit did not have an influence on the impact location of an individual electron hitting the screen, then in the above equation equality would hold. In the Copenhagen school this is interpreted as an exhibition of the wave-like nature of electrons, [^4] In textbooks one commonly encounters the description in which a wave function (on the detector) is associated to each slit and the pattern is said to emerge from an “interference term” [feynmanFeynmanLecturesPhysics1963]. However, such accounts are heuristic at best, since in QM $\sigma_{\text{double}}$ is obtained from the time evolution of a single initial wave function [kalbermannSingleDoubleslitScattering2002]. whereas the pointlike impacts are interpreted as an exhibition of their particle-like nature. Wave particle duality states that neither view provides a full picture, they are complementary.
@@
However, in the 1960s Landé expressed a different perspective on the matter: according to him [landeQuantumFactFiction1969], the electrons are pointlike particles and it is only their collective, statistical behavior that is wavelike.
@@
Indeed, the attentive reader may have noticed that the above argument does not rule out that electrons are particles, it only rules out certain notions of locality. For instance, an electron could pass through one of the slits, take a turn into the second slit and then go back through the first one before hitting the screen. Or, following de Broglie [bacciagaluppiQuantumTheoryCrossroads2009], one could imagine that the electron is pointlike, yet accompanied by a wave field in its immediate vicinity, which interacts with the other slit and in result changes the electron trajectory [couderSingleParticleDiffractionInterference2006, pucciWalkingDropletsInteracting2018]. At least qualitatively, both of these behaviors could explain Eq. (3.1). Of course, the argument I put forward here is that physically peculiar behavior can statistically lead to Eq. (3.1), it is not to commit to a particular explanation. In this sense, it seems difficult to evade the conclusion that Landé’s criticism of wave particle duality is justified.
@@
What Landé advocated for is a probabilistic understanding of quantum theory: while quantum theory does not predict the positions of the particles, it does predict the respective probabilities. Following Ballentine [ballentineStatisticalInterpretationQuantum1970, pechenkinStatisticalEnsembleInterpretation2022], quantum theory is concerned with ensembles of similarly prepared systems (of particles and radiation). An ensemble is a theoretical abstraction, consisting of an infinite number of samples. Each sample is a particular quantum system in the aforementioned sense. [^5] In the literature the distinction between the ensemble and the sample is often blurred. In particular, it depends on the context, whether the term ‘quantum system’ refers to a sample, an ensemble, or even a class of ensembles. Here I abide by this customary usage of the term. In line with a frequentist understanding of probability [richardvonmisesProbabilityStatisticsTruth1957] and staying on a heuristic level, the statistics of $n$ samples approach the respective probabilities of the ensemble as $n\to\infty$.
@@
In the above example this means that the relative impact frequencies of the electrons hitting the screen after $n$ runs will approach the probability density $\sigma_{\text{double}}$ as $n\to\infty$, appropriately understood. For $\sigma_{1}$ and $\sigma_{2}$ this is also the case, yet they refer to different physical contexts – i.e. different “similarly prepared systems” – and therefore belong to distinct ensembles.
@@
To summarize my argument: at least on a qualitative level, quantum phenomena like the double slit experiment do not conflict with Kolmogorov’s theory of probability.
@@
Rather, the applicability of Kolmogorov’s axioms to this physical domain is to be tested quantitatively. This can only be achieved indirectly by constructing and testing respective physical theories.
@@
Historically there have been many objections to such probabilistic approaches. [^6] See e.g. Refs. [vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955, kochenProblemHiddenVariables1975, puseyRealityQuantumState2012, dieksNeumannImpossibilityProof2017]. In Chap. 5 of [morettiFundamentalMathematicalStructures2019] Moretti has spelled out many of the underlying mathematical theorems. Addressing all of them would go beyond the scope of this article, so I shall focus on those which I deem most relevant in this context.
@@
First, if one follows the common interpretation of the uncertainty principle, then in quantum theory the notion of a point particle – let alone its trajectory – is not a sensible concept. The common interpretation is, however, not uncontested. Already in the 1930s Popper objected to it on the basis of a probabilistic view on quantum theory [popperZurKritikUngenauigkeitsrelationen1934]: [^7] See also §78 in [popperLogicScientificDiscovery1959]. Landé [landeQuantumFactFiction1969] and Fényes [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952] argued along similar lines. the assertion that quantum theory is concerned with probabilities implies that Heisenberg’s inequality cannot be interpreted as a statement on an individual particle. On the contrary, it ought to be understood as a probabilistic statement on the respective (quantum-mechanical) standard deviations (for the ensemble). That is, in the ensemble interpretation Heisenberg’s inequality is not a statement on the “localizability” of individual particles and can therefore not be employed as an argument against the particle model.

### v5

@@ Abstract
By formulating the axioms of quantum mechanics, von Neumann also laid the foundations of a “quantum probability theory”. As such, it is regarded a generalization of the “classical probability theory” due to Kolmogorov. Outside of quantum physics, however, Kolmogorov’s axioms enjoy universal applicability. This raises the question of whether quantum physics indeed requires such a generalization of our conception of probability or if von Neumann’s axiomatization of quantum mechanics was contingent on the absence of a general theory of probability in the 1920s.
This work argues in favor of the latter position. In particular, it shows how to construct a mathematically rigorous theory for non-relativistic $N$-body quantum systems subject to a time-independent scalar potential, which is based on Kolmogorov’s axioms and physically natural random variables. Though this theory is provably distinct from its quantum mechanical analog, it nonetheless reproduces central predictions of the latter. Further work may make an empirical comparison possible. Moreover, the approach can in principle be adapted to other classes of quantum-mechanical models.
Part II of this series discusses the empirical violation of Bell inequalities in the context of this approach. Part III addresses the projection postulate and the question of measurement.
@@ 1 Introduction
For almost a century now, the theory of quantum mechanics has stood firm at the core of our modern understanding of the physical microcosm. Still, as evinced by the multitude of interpretations in existence today [freirejuniorOxfordHandbookHistory2022], the theory continues to mystify even experts in the subject. This state of affairs has given quantum mechanics the idiosyncratic status of what one could call a well-established theory with contested foundations [freirejuniorQuantumDissidentsRebuilding2015, bacciagaluppiQuantumTheoryCrossroads2009].
@@
So far the following aspect of this problem has only received comparatively little attention [szaboCriticalReflectionsQuantum2001, accardiTopicsQuantumProbability1981, accardiNonKolmogorovProbability2009, fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952, nelsonDerivationSchrodingerEquation1966, nagasawaLocalityHiddenvariableTheories1997, redeiQuantumProbabilityTheory2007, streaterClassicalQuantumProbability2000, frohlichQuantumProbabilityTheory2015, reddigerMadelungPictureFoundation2017, reddigerProbabilisticFoundationNonrelativistic2022, reddigerMathematicalTheoryMadelung2023, mitaSchrodingersEquationDiffusion2021]: across the empirical sciences, quantum mechanics and its descendants are the only scientific theories that are not based on the mathematical theory of probability, as developed by Kolmogorov in 1933 [kolmogoroffGrundbegriffeWahrscheinlichkeitsrechnung1933]. Since von Neumann’s axiomatization of quantum mechanics [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927, vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955] predates Kolmogorov’s work, it therefore stands to reason, whether quantum physics indeed requires a different notion of probability or if von Neumann’s axiomatization of quantum mechanics was contingent on the absence of a general, mathematical theory of probability in the 1920s.
Should Kolmogorov’s theory of probability indeed be applicable to the description of quantum phenomena, then “quantum probability theory” would lose its physical support and contemporary quantum theory, both non-relativistic and relativistic, would require a thorough revision in terms of its underlying probability theory.
While the need for a “quantum probability theory” is widely defended in the literature [accardiTopicsQuantumProbability1981, accardiNonKolmogorovProbability2009, redeiQuantumProbabilityTheory2007, streaterClassicalQuantumProbability2000, frohlichQuantumProbabilityTheory2015], the counter-hypothesis has, so far, not been ruled out with certainty. The mere incompatibility of quantum mechanics with “classical probability theory”, as various so called “no go theorems” [vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955, kochenProblemHiddenVariables1975, puseyRealityQuantumState2012, dieksNeumannImpossibilityProof2017] aim to show, can hardly be taken serious as physical arguments (cf. [bellImpossiblePilotWave1982] and Chap. 1 in [reddigerProbabilisticFoundationNonrelativistic2022]), for those derive their relevance from the following implicit assertion: quantum mechanics is the only (non-relativistic) theory, that can correctly describe the data. On the contrary, it is an epistemological truism, that no amount of data can uniquely determine a single scientific theory. The scientific method demands that competing theories are empirically tested and compared on their own merits, if possible. That research in the history of the subject suggests the presence of a long-standing, systematic bias against theoretical approaches competing with (Copenhagen) quantum mechanics [cushingQuantumMechanicsHistorical1994, bellerQuantumDialogueMaking1999, freirejuniorQuantumDissidentsRebuilding2015], emphasizes the need for both scientific rigor and openness in this debate.
Accordingly, the overarching physical question of the applicability of Kolmogorov’s theory of probability to this domain of science can only be indirectly addressed by constructing specific competitor theories on this basis and following the scientific method.
@@
This work shows that for non-relativistic $N$-body quantum systems subject to a time-independent scalar potential, it is indeed possible to construct a theory based on Kolmogorov’s axioms and physically natural random variables, which reproduces central predictions of quantum mechanics (cf. Lem. 1, Thm. 1, and Prop. 1). Furthermore, the constructed theory is shown to be capable of making distinct predictions (cf. Ex. 3), so that the two theories can in principle be compared empirically.
@@
The main line of reasoning is as follows: First and with a historical focus, we consider the general relationship between mathematical probability theory and quantum mechanics in Sec. 2. Sec. 3 provides a probabilistic view on the double slit experiment and quantum theory as a whole, in line with the ensemble interpretation of quantum mechanics [ballentineStatisticalInterpretationQuantum1970, pechenkinStatisticalEnsembleInterpretation2022]. This physically motivates the implementation of Kolmogorov’s axioms in quantum theory via the Born rule for position (Sec. 4, in particular Lem. 1). Corresponding observables can be obtained from Madelung’s reformulation of the Schrödinger equation (Sec. 5), in line with [reddigerMadelungPictureFoundation2017]. This suggests that the theory, geometric quantum theory [reddigerMadelungPictureFoundation2017], can be erected upon the Madelung equations in conjunction with the Born rule. Yet a recent review [reddigerMathematicalTheoryMadelung2023] showed that the mathematical theory of the Madelung equations is still underdeveloped. Hence and at least for the time being, another approach is needed. Instead, one may develop a “hybrid theory”, in which the dynamics is borrowed from quantum mechanics and the basic quantities are rigorously defined in the context of said dynamics. For the aforementioned class of quantum systems this is done in Sec. 6. In result, an alternative theory for the aforementioned quantum system, which is based on Kolmogorov’s axioms, is indeed obtained.
The reader, who is primarily interested in the analytical results of this work, is invited to skip Secs. 2, 3, and 5, consulting Tab. 2 when needed. A synopsis of the theory developed here and a guideline for how to construct similar ones from other (classes of) quantum mechanical models is given in Sec. 7.
@@ 2 Probability and quantum mechanics
{tabu}
to X[1.4,l,m] X[3,l,m] X[3,l,m] von Neumann (pure state) Kolmogorov
Mathematical setting Separable, complex Hilbert space $\left(\mathcal{H},\left\langle\cdot\text{,}\cdot\right\rangle\right)$ Measure space $(\Lambda,\mathcal{A})$
Object that determines probabilities State $\Psi\in\mathcal{H}$ with $\sqrt{\left\langle\Psi\text{,}\Psi\right\rangle}=\lVert\Psi\rVert=1$ Measure $\mathbb{P}\colon\mathcal{A}\to[0,1]\ ,\ A\mapsto\mathbb{P}(A)$  with $\mathbb{P}(\Lambda)=1$
“Observable”/ random variable Unbounded self-adjoint operator $\hat{X}\colon\operatorname{dom}\hat{X}\to\mathcal{H}$       with dense domain $\operatorname{dom}\hat{X}$ in $\mathcal{H}$ Measurable function $X\colon\Lambda\to\mathbb{R}$
Probability to obtain values in $I\in\mathcal{B}(\mathbb{R})$ $\frac{\left\lvert\left\langle\Psi\text{,}P_{I}\Psi\right\rangle\right\rvert^{2}}{\lVert P_{I}\Psi\rVert^{2}}=\int_{I}\mathrm{d}\negthinspace\left\langle\Psi\text{,}P_{\cdot}\Psi\right\rangle$ $\int_{X^{-1}(I)}\mathrm{d}\mathbb{P}=\int_{I}\mathrm{d}\bigl{(}\mathbb{P}\circ X^{-1}\bigr{)}$
[…]
Table 1: Direct comparison of the two probability theories. $\mathcal{B}(\mathbb{R})$ denotes the Borel sets on $\mathbb{R}$. The map $I\mapsto P_{I}$ is the projection valued measure for the operator $\hat{X}$ and, accordingly, $I\mapsto\left\langle\Psi\text{,}P_{I}\Psi\right\rangle$ is the spectral measure (cf. Sec. VIII.3 in [reedMethodsModernMathematical1972]). Note the close analogy between the spectral measure and $\mathbb{P}\circ X^{-1}$.
For an introduction to mathematical probability theory the reader is referred to [klenkeProbabilityTheoryComprehensive2013]. Operator theory with a focus on quantum mechanics is treated, for instance, in [morettiFundamentalMathematicalStructures2019].
@@
The two modern mathematical theories of probability were developed by the Hungarian mathematician John von Neumann (1903–1957) [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927, vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955] and the Russian mathematician Andrej N. Kolmogorov (1903–1987) [kolmogoroffGrundbegriffeWahrscheinlichkeitsrechnung1933]. That today there are two such theories, points at a potential discrepancy in the modern foundations of probability theory (as a structural theory of nature).
@@
Historically, it is known that both researchers knew each other and followed each other’s work [monastyrskiiJohnNeumann2004]. Still, the author did not find any documented evidence of a discussion among the two about this particular topic, even after searching in the literature and a consultation of a monograph on the history of modern probability theory [vonplatoCreatingModernProbability1994]. On the contrary, Parthasarathy [parthasarathyObituaryAndreiNikolaevich1988] writes that Kolmogorov showed little interest in quantum mechanics, adding that “it is a blessing to have enough ground left for future mathematicians to till.”
To the author’s knowledge, the Hungarian physicist Fényes was the first person in the literature to claim the validity of Kolmogorov’s axioms in the quantum domain [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952]. The physical need for a “quantum probability theory” was explicitly questioned by Szabó [szaboCriticalReflectionsQuantum2001], though on the basis of different arguments and following a different approach than the one pursued here.
@@
A direct comparison of the two probability theories for pure states is given in Tab. 2. In the contemporary literature this conflict in the foundations of probability is resolved by showing that von Neumann’s “quantum probability theory” may be viewed as a “non-commutative” generalization of the “classical probability theory” discovered by Kolmogorov [accardiTopicsQuantumProbability1981, streaterClassicalQuantumProbability2000, redeiQuantumProbabilityTheory2007, szaboCriticalReflectionsQuantum2001].
@@
While this accords with the general philosophy that quantum mechanics is to be viewed as a generalization of “classical mechanics”, the particular historical context of the development of the two theories may be put forward as an argument against this conventional point of view.
@@
Mostly guided by the ideas of Heisenberg, Dirac, Born and Jordan, von Neumann developed his theory of probability with the explicit goal of axiomatizing quantum mechanics [vonneumannMathematischeBegrundungQuantenmechanik1927, vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927]. That, and not the development of a probability theory, was his aim. Indeed, in his most important article on the topic [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927] von Neumann noted that he was merely formalizing “the common calculus of probability.” Clearly, he was not aware that one day his notion of probability would be in conflict with what is generally associated with that term.
@@
Kolmogorov, on the other hand, took a direct approach to the question via measure theory. His fundamental work [kolmogoroffGrundbegriffeWahrscheinlichkeitsrechnung1933] was published in 1933. This is one year after von Neumann’s famous book on the mathematical foundations of quantum mechanics [vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955] was published and a total of six years after von Neumann’s article on the topic [vonneumannWahrscheinlichkeitstheoretischerAufbauQuantenmechanik1927] appeared.
@@
One should add to this, that today Kolmogorov’s theory of probability is successfully employed in virtually all of the sciences, except quantum theory. Convincing examples outside of quantum physics, where von Neumann’s probability theory is required, seem to be missing (see [accardiNonKolmogorovProbability2009] for a different view).
Therefore, the hypotheses put forward here are that, one, von Neumann’s axiomatization was contingent on the absence of a general, mathematical theory of probability at the time and, two, that the physically correct theory of probability is indeed the one discovered by Kolmogorov.
If this is the case, then contemporary quantum theory, both non-relativistic and relativistic, requires a thorough revision in terms of its underlying probability theory.
In order to put these hypotheses to scientific scrutiny, the question of whether an empirically successful quantum theory can be constructed on the basis of Kolmogorov’s axioms is to be investigated. What follows shows that this may indeed be possible.
@@ 3 Physical considerations
Before addressing any mathematical details, the reader shall be briefly introduced to the central physical ideas.
For this purpose, we begin with a discussion of the double slit experiment with electrons. The argument in principle, also applies to other material particles; on the topic of electromagnetic radiation the reader is referred to Planck’s remark in [einsteinUeberEntwicklungUnserer1969]. The character of the double slit experiment as an “examplar” in the sense of Kuhn [kuhnStructureScientificRevolutions2012] was also recognized by Feynman, who wrote that “it has in it the heart of quantum mechanics” and that “[i]n reality, it contains the only mystery” [feynmanFeynmanLecturesPhysics1963].
For the sake of argument, we shall assume that in the experiment only a single electron is released at a time, so that one can view the runs as mutually independent. Then, if an electron impinges on the detector screen, it creates a single, pointlike impact. Yet, as one repeats the experiment a large number of times, an interference pattern begins to emerge on the screen.
In textbook accounts, wave functions are commonly associated with each slit and the pattern is said to emerge from an “interference term”, resulting out of the absolute modulus squared of the superposition [feynmanFeynmanLecturesPhysics1963]. However, such accounts are heuristic and somewhat misleading, since in quantum mechanics the pattern must be viewed as a consequence of the time evolution of a single initial wave function in the presence of the double slit [kalbermannSingleDoubleslitScattering2002].
Still, the argument, that this pattern cannot result from particles in the usual sense of the word [feynmanFeynmanLecturesPhysics1963], needs to be addressed. The argument is that the pattern is not the same as the conjunction of the two patterns one obtains with one of the slits closed. That is, if $\sigma_{\text{double}}$ is the ($2$-dimensional, time-independent) probability density of the impact locations on the detector screen for the double slit experiment and $\sigma_{i}$ with $i\in\{1,2\}$ are the probability densities for the experiment with one of the slits closed, then we have
@@
Indeed, if the presence of the other slit did not have an influence on the impact location of an individual electron hitting the screen, then in the above equation equality would hold. In the Copenhagen school this is interpreted as an exhibition of the wave-like nature of electrons, whereas the pointlike impacts are interpreted as an exhibition of their particle-like nature. Wave particle duality states that neither view provides a full picture, they are complementary.
@@
In the 1960s Landé expressed a different perspective on the matter: according to him [landeQuantumFactFiction1969], the electrons are pointlike particles and it is only their collective, statistical behavior that is wavelike.
@@
Indeed, the attentive reader may have noticed that the above argument does not rule out that electrons are particles, it only rules out naive notions of locality. For instance, an electron could pass through one of the slits, take a turn into the second slit and then go back through the first one before hitting the screen. Or, following de Broglie [bacciagaluppiQuantumTheoryCrossroads2009], one could imagine that the electron is pointlike, yet accompanied by a wave field in its immediate vicinity, which interacts with the other slit and in result changes the electron trajectory [couderSingleParticleDiffractionInterference2006, bushQuantumMechanicsWrit2010, pucciWalkingDropletsInteracting2018]. Such trajectories must differ, when one of the slits is closed. Hence, both of these behaviors could explain Eq. (3.1), at least qualitatively. Of course, the argument put forward here is that physically peculiar behavior can statistically lead to Eq. (3.1), it is not to commit to a particular explanation. In this sense, it seems difficult to evade the conclusion that Landé’s criticism of wave particle duality is indeed justified.
@@
What Landé advocated for is a probabilistic understanding of quantum theory: while quantum theory does not predict the positions of the particles, it does predict the respective probabilities. Following Ballentine [ballentineStatisticalInterpretationQuantum1970, pechenkinStatisticalEnsembleInterpretation2022], quantum theory is concerned with ensembles of similarly prepared systems (of particles and radiation). An ensemble is a theoretical abstraction, consisting of an infinite number of samples. Each sample is a particular quantum system in the aforementioned sense (though we shall not be overly pedantic in our usage of the term “quantum system” here). In line with a frequentist understanding of probability [richardvonmisesProbabilityStatisticsTruth1957] and staying on a heuristic level, the statistics of $n$ samples approach the respective probabilities of the ensemble as $n\to\infty$, as long as the random variables in question are mutually independent and identically distributed for each run.
@@
In the above example this means that the relative impact frequencies of the electrons hitting the screen after $n$ runs will approach the probability density $\sigma_{\text{double}}$ as $n\to\infty$, appropriately understood. For $\sigma_{1}$ and $\sigma_{2}$ this is also the case, yet they refer to different physical contexts, different “similarly prepared systems” that is, and therefore belong to distinct ensembles.
@@
To summarize the argument: the results of the double slit experiment do not conflict with Kolmogorov’s theory of probability, at least on a qualitative level. The general claim is that this is also true for other quantum phenomena.
@@
Rather, the applicability of Kolmogorov’s axioms to this physical domain is to be tested quantitatively. Again, this can only be achieved indirectly by constructing and testing respective physical theories.
@@
Historically, there have been many objections to such probabilistic approaches. The interested reader is referred, for instance, to [vonneumannMathematischeGrundlagenQuantenmechanik1932, vonneumannMathematicalFoundationsQuantum1955, kochenProblemHiddenVariables1975, puseyRealityQuantumState2012, dieksNeumannImpossibilityProof2017], as well as Chap. 5 in Moretti’s book [morettiFundamentalMathematicalStructures2019]. Addressing all of these objections individually would go beyond the scope of this article, so we focus on those, which are arguably the strongest in this context.
@@
First, if one follows the common interpretation of the uncertainty principle, then in quantum theory the notion of a point particle, let alone its trajectory, is not a sensible concept. The common interpretation is, however, not uncontested. Already in the 1930s Popper objected to it on the basis of a probabilistic view on quantum theory (see [popperZurKritikUngenauigkeitsrelationen1934], §78 in [popperLogicScientificDiscovery1959], and [shieldsHistoricalSurveySir2012] for a historical account): the assertion that quantum theory is concerned with probabilities implies that Heisenberg’s inequality cannot be interpreted as a statement on an individual particle. On the contrary, it ought to be understood as a probabilistic statement on the respective (quantum-mechanical) standard deviations (for the ensemble). That is, in the ensemble interpretation Heisenberg’s inequality is not a statement on the “localizability” of individual particles and can therefore not be employed as an argument against the particle model. Landé [landeQuantumFactFiction1969] and Fényes [fenyesWahrscheinlichkeitstheoretischeBegruendungUnd1952] argued along similar lines.